# TopFormer: Topology-Aware Authorship Attribution of Deepfake Texts with Diverse Writing Styles

### v2

@@ Abstract
Recent advances in Large Language Models (LLMs) have enabled the generation of open-ended high-quality texts, that are non-trivial to distinguish from human-written texts. We refer to such LLM-generated texts as deepfake texts. There are currently over 72K text generation models in the huggingface model repo. As such, users with malicious intent can easily use these open-sourced LLMs to generate harmful texts and dis/misinformation at scale. To mitigate this problem, a computational method to determine if a given text is a deepfake text or not is desired–i.e., Turing Test (TT). In particular, in this work, we investigate the more general version of the problem, known as Authorship Attribution (AA), in a multi-class setting–i.e., not only determining if a given text is a deepfake text or not but also being able to pinpoint which LLM is the author. We propose TopFormer to improve existing AA solutions by capturing more linguistic patterns in deepfake texts by including a Topological Data Analysis (TDA) layer in the Transformer-based model. We show the benefits of having a TDA layer when dealing with imbalanced, and multi-style datasets, by extracting TDA features from the reshaped $pooled\_output$ of our backbone as input. This Transformer-based model captures contextual representations (i.e., semantic and syntactic linguistic features), while TDA captures the shape and structure of data (i.e., linguistic structures). Finally, TopFormer, outperforms all baselines in all 3 datasets, achieving up to 7% increase in Macro F1 score.
@@ 1 Introduction
The first step to mitigate these limitations of LLMs starts with our ability to determine whether a text is generated by a particular LLM (deepfake text) or not. This task is known as Turing Test (TT) [40]. Further, in this work, generalizing the TT problem further, we are interested in not only determining if a text is a deepfake text or not but also pinpointing which LLM is the author, known as the Authorship Attribution (AA) problem [41]. See the illustration of AA in Figure 1. The AA problem in a multi-class setting has not been as rigorously studied as the TT problem has. Naturally, AA is substantially more challenging than TT. However, with the ever-increasing number of popular LLMs that people can use, we believe that it is no longer sufficient to just ask the binary question of “is this written by human or machine?" Solving the AA problem enables more fine-grained and targeted defense tools for users and platforms (e.g., a detector specifically trained and designed to detect ChatGPT deepfake texts). AA solutions can also help researchers and policymakers understand the capacity and limitations of different LLMs, and study which LLMs are more vulnerable to misuse and abuse in what context (e.g., political propaganda, terrorism recruitment, phishing).
@@ 2.1 Authorship Attribution of Deepfake Texts
Thus, for the stylometric solution, several solutions are using POS tags, lexical dictionaries, etc. to capture an author’s unique writing style [39, 18]. Next for the deep learning solutions, researchers usually fine-tune BERT models and BERT variants [40, 14, 17]. However, as fine-tuning can be computationally expensive and requires a lot of data which does not always exist, some researchers have proposed an unsupervised technique - statistical-based solutions [10, 25]. However, while deep learning-based classifiers perform very well, they are highly susceptible to adversarial perturbations [41]. Therefore, researchers propose hybrid-based solutions that fuse both deep learning and statistical-based or stylometric solutions to improve performance and robustness [20, 4]. More recently, prompt-based techniques are being used for attribution of deepfakes [16, 19]. Lastly, for the human-based approaches, to improve human detection of deepfake texts, researchers have utilized two main techniques - training [10, 42] and not training [14].
@@ 4 TopFormer: Topology-Aware Attributor
To build this TDA-infused Transformer-based model, we focus on the four layers needed to convert the vanilla Transformer-based (RoBERTa) model to a Topological Transformer-based model (TopFormer) - (1) pre-trained weights of the backbone model, (2) dropout layer with probability $p$=0.3, (3) Topological layer for calculating, and (4) Linear transformation layer. See Figure 2 for a flow chart describing the architecture of TopFormer with the 4 layers.

### v3

@@ Abstract
Recent advances in Large Language Models (LLMs) have enabled the generation of open-ended high-quality texts, that are non-trivial to distinguish from human-written texts. We refer to such LLM-generated texts as deepfake texts. There are currently over 72K text generation models in the huggingface model repo. As such, users with malicious intent can easily use these open-sourced LLMs to generate harmful texts and dis/misinformation at scale. To mitigate this problem, a computational method to determine if a given text is a deepfake text or not is desired–i.e., Turing Test (TT). In particular, in this work, we investigate the more general version of the problem, known as Authorship Attribution (AA), in a multi-class setting–i.e., not only determining if a given text is a deepfake text or not but also being able to pinpoint which LLM is the author. We propose TopFormer to improve existing AA solutions by capturing more linguistic patterns in deepfake texts by including a Topological Data Analysis (TDA) layer in the Transformer-based model. We show the benefits of having a TDA layer when dealing with imbalanced, and multi-style datasets, by extracting TDA features from the reshaped $pooled\_output$ of our backbone as input. This Transformer-based model captures contextual representations (i.e., semantic and syntactic linguistic features), while TDA captures the shape and structure of data (i.e., linguistic structures). Finally, TopFormer, outperforms all baselines in all 3 datasets, achieving up to 7% increase in Macro F1 score. Our code and datasets are available at: https://github.com/AdaUchendu/topformer
@@ 1 Introduction
The first step to mitigate these limitations of LLMs starts with our ability to determine whether a text is generated by a particular LLM (deepfake text [^1] We call these LLM-generated texts, deepfake texts; However, there are other terms which include - AI-generated, auto-generated, machine-generated, synthetic, artificial, computer-generated, neural, and LLM-generated texts) or not. This task is known as Turing Test (TT) [41]. Further, in this work, generalizing the TT problem further, we are interested in not only determining if a text is a deepfake text or not but also pinpointing which LLM is the author, known as the Authorship Attribution (AA) problem [42]. See the illustration of AA in Figure 1. The AA problem in a multi-class setting has not been as rigorously studied as the TT problem has. Naturally, AA is substantially more challenging than TT. However, with the ever-increasing number of popular LLMs that people can use, we believe that it is no longer sufficient to just ask the binary question of “is this written by human or machine?" Solving the AA problem enables more fine-grained and targeted defense tools for users and platforms (e.g., a detector specifically trained and designed to detect ChatGPT deepfake texts). AA solutions can also help researchers and policymakers understand the capacity and limitations of different LLMs, and study which LLMs are more vulnerable to misuse and abuse in what context (e.g., political propaganda, terrorism recruitment, phishing).
@@ 2.1 Authorship Attribution of Deepfake Texts
Thus, for the stylometric solution, several solutions are using POS tags, lexical dictionaries, etc. to capture an author’s unique writing style [40, 18]. Next for the deep learning solutions, researchers usually fine-tune BERT models and BERT variants [41, 14, 17, 35, 1]. However, as fine-tuning can be computationally expensive and requires a lot of data which does not always exist, some researchers have proposed an unsupervised technique - statistical-based solutions [10, 25]. However, while deep learning-based classifiers perform very well, they are highly susceptible to adversarial perturbations [42]. Therefore, researchers propose hybrid-based solutions that fuse both deep learning and statistical-based or stylometric solutions to improve performance and robustness [20, 5]. More recently, prompt-based techniques are being used for attribution of deepfakes [16, 19, 43]. Lastly, for the human-based approaches, to improve human detection of deepfake texts, researchers have utilized two main techniques - training [10, 43] and not training [14].
@@ 3 Topological Data Analysis (TDA) Features
Algorithm 1 TopFormer Algorithm (Forward Pass)
1:  Input: Document $x$, pretrained transformer model $M(\cdot)$, fully-connected neural network $f(\cdot)$.
2:  Output: Predicted authorship label $y^{\prime}$
[…]
8:  Return: $y^{\prime}$
@@ 4 TopFormer: Topology-Aware Attributor
To build this TDA-infused Transformer-based model, we focus on the four layers needed to convert the vanilla Transformer-based (RoBERTa) model to a Topological Transformer-based model (TopFormer) - (1) pre-trained weights of the backbone model, (2) dropout layer with probability $p$=0.3, (3) Topological layer for calculating, and (4) Linear transformation layer. See Figure 2 for a flow chart describing the architecture of TopFormer with the 4 layers and Algorithm 1 for the pseudocode for (TopFormer).
@@ 6 Results
In addition, adding a TDA layer to RoBERTa increases the training time and inference time by about 0.5-2 hours, depending on data size. However, just inference time remains fast.
@@ 11 Ethical Statement
Acknowledgement
This work was in part supported by NSF awards #1820609, #2114824, and #2131144. Adaku thanks Dr. Bastian Rieck for the insightful discussions on how to improve our TDA technique. Also, Adaku would like to thank Dr. Charlie Dagli for reading the paper drafts and providing invaluable recommendations.